A 6565 kg rocket is being launched up. It is at rest on the launch pad when the rocket engine starts and supplies a thrust of 125,000 N. How fast will the rocket be traveling after 25 sec?
1
Expert's answer
2020-11-16T07:40:09-0500

According to the Newton's second law, the acceleration "a" of the rocket is:


"ma = F-mg\\
a = \dfrac{F}{m} -g"

where "m = 6565 kg", "F = 125000N" and "g = 9.81m\/s^2" is the gravitational acceleration.

The distance traveled under the constant acceleration during the time "t = 25s" is:


"d = \dfrac{at^2}{2} = \dfrac{t^2}{2}\left (\dfrac{F}{m} -g\right)\\
d = \dfrac{25^2}{2}\left (\dfrac{125000}{6565} -9.81\right)\approx 2884.5 m"

Answer. 2884.5 m.